# Fairness in Ranking under Disparate Uncertainty

### v3

@@ Appendix C Proofs
Now using $\frac{p_{i}}{nRel(A)}\leq\frac{\sigma^{PRP,A}[1]}{nRel(A)}$, and Equation (27) above,
@@
From Equation (32) and (33), $-\delta_{max}\leq\delta(\sigma_{k})\leq 0$ and thus EOR algorithm will add all the remaining elements from group A resulting in $\left|\delta(\sigma_{k})\right|\leq\delta_{max}$. Analogously, it can be shown that if all the elements from group A had been added by prefix $k$, adding the next element from group B would satisfy $\left|\delta(\sigma_{k})\right|\leq\delta_{max}$.
@@
When $\delta(\sigma^{EOR}_{k})=0$, by the definition of EOR criterion, we have that $\frac{nRel(A|\sigma^{EOR}_{k})}{nRel(A)}=\frac{nRel(B|\sigma^{EOR}_{k})}{nRel(B)}$. As a result, the total cost ($1-\frac{nRel(A|\sigma^{EOR}_{k})+nRel(B|\sigma^{EOR}_{k})}{nRel(A)+nRel(B)}$) as well as subgroup cost would be equal to
@@
Adding Equations (35), (36) and using (34), we get that
@@
This and Equation (34) are sufficient to claim that the total cost and subgroup costs of uniform policy given by $1-\frac{k}{n}$ will always be higher than the total cost and subgroup costs given by EOR ranking when $\delta(\sigma^{EOR}_{k})=0$. ∎
C.4. Proof for Proposition 5.4
See 5.4
Proof.
Let there be two rankings denoted by $\sigma_{k}^{EOR}$ and $\sigma_{k}^{PRP}$ for EOR and PRP respectively. We assume that $\sigma_{k}^{EOR}$ selects $k_{A}$ and $k_{B}$ elements from group A and B respectively. Similarly $\sigma_{k}^{PRP}$ selects $k^{\prime}_{A}$ and $k^{\prime}_{B}$ elements from group A and B respectively. Then $k_{A}+k_{B}=k^{\prime}_{A}+k^{\prime}_{B}=k$ where $k$ are the total top-k elements selected by each ranking.
Within each group both PRP and EOR select elements in decreasing order of relevances. If we assume that $nRel(A|\sigma_{k}^{PRP})\geq nRel(A|\sigma_{k}^{EOR})$, i.e that PRP selects more candidates from group A compared to EOR then $k^{\prime}_{A}>k_{A}$ and EOR selects more candidates from group B compared to PRP with $k_{B}>k^{\prime}_{B}$. We also have $p_{k_{B}}<p_{k^{\prime}_{B}}$ since within group ranking is presevered by both PRP and EOR.
Further, we assume that $\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR}$ is the set of elements selected by both $\sigma_{k}^{PRP}$ and $\sigma_{k}^{EOR}$ and includes elements up to $k_{A}$ in group A and $k^{\prime}_{B}$ in group B. EOR criterion for the rankings is as follows
$\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})$ $=$ $\frac{\sum_{i=1}^{k_{A}}p_{i}}{nRel(A)}-\frac{\sum_{j=1}^{k^{\prime}_{B}}p_{j}}{nRel(B)}$
$\delta(\sigma_{k}^{EOR})$ $=$ $\frac{\sum_{i=1}^{k_{A}}p_{i}}{nRel(A)}-\frac{\sum_{j=1}^{k_{B}}p_{j}}{nRel(B)}=\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})-\frac{\sum_{j=k^{\prime}_{B}+1}^{k_{B}}p_{j}}{nRel(B)}$
$\delta(\sigma_{k}^{PRP})$ $=$ $\frac{\sum_{i=1}^{k^{\prime}_{A}}p_{i}}{nRel(A)}-\frac{\sum_{j=1}^{k^{\prime}_{B}}p_{j}}{nRel(B)}=\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})+\frac{\sum_{i=k_{A}+1}^{k^{\prime}_{A}}p_{i}}{nRel(A)}$
where $\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})$ is the EOR criterion for the common elements selected by $\sigma^{PRP}$ and $\sigma^{EOR}$.
As a result, the cost differential between the two rankings is given by
Cost Differential $=$ $c(\text{Principal}|\sigma_{k}^{EOR})-c(\text{Principal}|\sigma_{k}^{PRP})$
$=$ $\frac{1}{nRel(A)+nRel(B)}[\sum_{i=k_{A}+1}^{k^{\prime}_{A}}p_{i}-\sum_{j=k^{\prime}_{B}+1}^{k_{B}}p_{j}]$
$=$ $\frac{1}{nRel(A)+nRel(B)}[nRel(A)(\delta(\sigma_{k}^{PRP})-\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR}))+nRel(B)(\delta(\sigma_{k}^{EOR})-\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR}))]$
$=$ $\frac{nRel(A)}{nRel(A)+nRel(B)}\delta(\sigma_{k}^{PRP})+\frac{nRel(B)}{nRel(A)+nRel(B)}\delta(\sigma_{k}^{EOR})-\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})$
Similarly, if $nRel(B|\sigma_{k}^{PRP})\geq nRel(B|\sigma_{k}^{EOR})$, i.e that PRP selects more candidates from group B compared to EOR, then the cost differential is given by
Cost Differential $=$ $c(\text{Principal}|\sigma_{k}^{EOR})-c(\text{Principal}|\sigma_{k}^{PRP})$
$=$ $\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})-\frac{nRel(B)}{nRel(A)+nRel(B)}\delta(\sigma_{k}^{PRP})-\frac{nRel(A)}{nRel(A)+nRel(B)}\delta(\sigma_{k}^{EOR})$
@@ Appendix D Extension to Multiple Groups $G$
The primal of the LP to find a solution $X$ for this problem is as follows
Objective
$\text{Maximize}_{x\geq 0}f(X)=\frac{P^{T}X}{\sum_{g=1}^{G}nRel(g)}$
subject to
(37) $X$$\leq$ 11 11
@@
Dual
$\text{Minimize}_{\lambda\geq 0}g(\lambda)=\delta(\sigma^{EOR}_{k})\overbrace{\sum_{\{A,B\}}(\lambda_{A,B}+\lambda_{B,A})}^{G(G-1)/2}+k\lambda_{k}+\sum_{i=1}^{n}\lambda_{i}^{\prime}$
subject to
(39) $\sum_{\{A,B\}}Q_{A,B}(\lambda_{A,B}-\lambda_{B,A})+\lambda_{k}+\lambda^{\prime}\geq\frac{P}{\sum_{g}nRel(g)}$
@@
We construct $\lambda^{\prime}_{i}$ corresponding to constraint (37) and $\lambda_{k}$ corresponding to constraint (38) below.
@@
$\phi=\sum_{g=1}^{G}nRel(g)\left(\sum_{\{A,B\}}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|\right)$ and $\delta_{max}=\max_{g}\left\{\frac{\sigma^{PRP,g}[1]}{nRel(g)}\right\}$ from Lemma D.4. ∎
@@
This lemma follows directly from the EOR ranking principle of choosing the candidate that minimizes $\delta(\sigma^{EOR}_{k})$ defined according to Equation  (8).
@@
For every pair of $\lambda_{A,g}$ and $\lambda_{g,A}$, where $g\in\{1,\cdots G\}$ and $g\neq A$, only one of $\lambda_{A,g},\lambda_{g,A}$ is $\geq 0$. Each of the $G-1$ terms inside the summation in Equation (44) reduces to the two group case as follows. For $i>k_{A}$ and each {$A,g$}, the term evaluates to $\leq 0$ using Lemma C.2 and thus $\lambda_{i}^{\prime}$ is clipped to 00. Similarly, for $i\leq k_{A}$ and each {$A,g$} the term evaluates to $\geq 0$ and thus $\lambda_{i}^{\prime}\geq 0$.
@@
We now show that $\lambda_{k}\geq 0$. From Equation (42),
@@
Each of the $G-1$ terms inside the summation in Equation (45) reduces to the two group case. For each {$A,g$}, the term evaluates to $\geq 0$ using Lemma C.2 and thus $\lambda_{k}\geq 0$.
@@
Using the fact that $\lambda_{i}^{\prime}=0$ for $i>k_{A}$ from Lemma D.2, and substituting $\lambda_{k}$ from Equation (42), we get
@@
Each of the $G-1$ terms inside the summation in Equation (47) reduces to the two group case. For each {$A,g$}, the term evaluates to $\frac{p_{i}}{(G-1)\sum_{g}nRel(g)}$ using Lemma C.3 and thus the corresponding duality constraint is satisfied.
@@
Using the fact that $\lambda_{i}^{\prime}\geq 0$ for $i\leq k_{A}$ from Lemma D.2, and substituting $\lambda_{k}$ from Equation (42), $\lambda_{i}^{\prime}$ for $i\leq k_{A}$ in (43), we get
@@
We argue by an inductive argument similar to the proof of Theorem 5.2. Consider the base case of $k=1$, when the first element is to be selected. The EOR algorithm will select according to Equation (8) resulting in the lower $\delta(\sigma_{k=1})$. Thus, $\delta(\sigma_{k=1})$ is clearly $\leq\delta_{max}$.
@@
At the next prefix $k$, if the group that is selected has $\frac{nRel(.|\sigma_{k})}{nRel(.)}\leq\frac{nRel(g|\sigma_{k-1})}{nRel(g)}$, then $\delta(\sigma_{k})\leq\delta_{max}$. Note that $\delta(\sigma_{k})$ is always non-negative by definition from Equation (8).
@@
Equation (D.1) holds since group $1$ is now the group with maximum relevance proportion after adding $p_{i}$ the top most current element from group $1$. Group $2$ becomes the group with minimum relevance proportion.
@@
We have shown above that if the group with lowest relevance proportion at prefix $k-1$ (group $1$ in this case) is selected and its relevance proportion now exceeds the group with the highest relevance proportion at prefix $k-1$ (group $g$ in the case above), then $\delta(\sigma_{k})\leq\delta_{max}$. Thus, we can say that at least one group exists that satisfies $\delta_{max}$ EOR constraint at prefix $k$.
This completes the proof that the EOR algorithm always provides a feasible ranking that satisfies
$\delta_{max}=\max_{g\in\{1\cdots G\}}\left\{\frac{\sigma^{PRP,g}[1]}{nRel(g)}\right\}$ for $G$ groups. ∎
@@ E.1. Baselines
Thompson Sampling Ranking Policy ($\pi^{TS}$)
(Singh et al., 2021) For $\pi^{TS}$, binary relevances are drawn according to $r_{i}\sim\mathbb{P}(r_{i}|\mathcal{D})$, and candidates are sorted in decreasing order of relevance $r_{i}$ with their ranking randomized for the same value of relevance $r_{i}$.
@@
Demographic Parity ($\pi^{DP}$
) Candidates in each group are sorted in decreasing order of $\mathbb{P}(r_{i}|\mathcal{D})$ and selected such that the following constraint is minimized. This constraint (51) is similar to the different statistical parity variations introduced in (Yang and Stoyanovich, 2016).
@@
where $S(.)$ represents the size of the group. For a fair comparison with $\sigma^{EOR}$, we use Algorithm 1 and instead of minimizing Equation 6, we minimize the above demographic parity constraint (51). We now discuss other variations of proportional representation constraints that have been introduced in prior literature (Celis et al., 2017, 2020, 2021). Generally, these constraints require that the disadvantaged group selected is at least a specific proportion $\alpha$ of top k.
@@
Equation (52) is used as the fairness constraint while maximizing the utility to the principal. This type of representational constraint by definition requires the designation of a disadvantaged group. By designating B as the disadvantaged group, the constraint for proportional Rooney-Rule policy (Shen et al., 2023), which we denote by $\pi^{PRR}$ is given by $\alpha=\frac{S(B)}{S(A)+S(B)}$
@@
Figure 9. EOR fairness $\delta(\sigma_{k})$, costs of the ranking policies, and DCG Utility for Synthetic dataset with proportional Rooney-Rule like constraint, $\pi^{PRR}$. For group A we draw $S(A)=30$ relevance probabilities from $Powerlaw(\eta=5)$, and then draw for group B from $Powerlaw(\eta=0.5)$ until $nRel(A)\approx nRel(B)$.
@@
This criterion is anchored on the principle that a top-k ranking is fair when the proportion of disadvantaged candidates selected doesn’t fall far below a required minimum proportion $p$. This is formalized with a Binomial distribution, and a confidence level ($1-\alpha$). A function of the binomial cdf is computed apriori and is used as an input in the FAIR Algorithm. Since Binomial(p=0.5,n) corresponds to a ranking where at each position, a candidate from either group is selected randomly, FAIR is a ”softened” version of demographic parity (DP). As a result, FAIR is fundamentally different from Axiom 1 and 4.1 derived from the uniform lottery fairness because, unlike DP, the uniform lottery is anchored on selecting an equal fraction of relevance from each group. Unlike $\pi^{EOR}$, $\pi^{FS}$ is oblivious to the relevance distribution and thus cannot take disparate uncertainty into account. FAIR also requires the normative designation of a disadvantaged group.
@@
For $\pi^{RA}$, we modify the baseline for fair rank aggregation in (Cachel and Rundensteiner, 2023) as follows. In fair rank aggregation, all $n$ candidates are ranked by $m$ voters to achieve a ranking with maximum consensus accuracy, where consensus may be according to different aggregation methods while achieving fairness of exposure w.r.t groups. (Cachel and Rundensteiner, 2023) proposes an algorithm that finds the consensus maximizing ranking and then swaps the candidates such that the equality of exposure is satisfied in that ranking. To adapt this baseline, we use the ranking from utility maximizing $\pi^{PRP}$ as the consensus ranking and use their algorithm to swap elements in PRP ranking until the exposure constraint below is satisfied,
@@
A threshold of $0.95$ is used in experiments and on average over 100 runs, an exposure of $0.96\pm{0.01},0.96\pm{0.00},0.97\pm{0.00}$ is achieved for high, medium, and low levels of disparate uncertainty respectively.

### v4

@@ Appendix C Proofs
Now using $\frac{p_{i}}{nRel(A)}\leq\frac{\sigma^{PRP,A}[1]}{nRel(A)}$, and Eq. (26) above,
@@
From Eqs. (31) and (32), $-\delta_{max}\leq\delta(\sigma_{k})\leq 0$ and thus EOR algorithm will add all the remaining elements from group A resulting in $\left|\delta(\sigma_{k})\right|\leq\delta_{max}$. Analogously, it can be shown that if all the elements from group A had been added by prefix $k$, adding the next element from group B would satisfy $\left|\delta(\sigma_{k})\right|\leq\delta_{max}$.
@@
When $\delta(\sigma^{EOR}_{k})=0$, by the definition of EOR fairness, we have that $\frac{nRel(A|\sigma^{EOR}_{k})}{nRel(A)}=\frac{nRel(B|\sigma^{EOR}_{k})}{nRel(B)}$. As a result, the total cost ($1-\frac{nRel(A|\sigma^{EOR}_{k})+nRel(B|\sigma^{EOR}_{k})}{nRel(A)+nRel(B)}$) as well as subgroup cost would be equal to
@@
Adding Eqs. (34), (35) and using (33), we get that
@@
This and Eq. (33) are sufficient to claim that the total cost and subgroup costs of uniform policy given by $1-\frac{k}{n}$ will always be higher than the total cost and subgroup costs given by EOR ranking when $\delta(\sigma^{EOR}_{k})=0$. ∎
@@ Appendix D Extension to Multiple Groups $G$
The LP to find a solution $X$ for this problem is formulated as follows
(Primal) $\max_{x\geq 0}\quad f(x)$ $=\frac{P^{T}X}{\sum_{g=1}^{G}nRel(g)}$
@@
$G(G-1)$ constraints $\left\{\begin{array}[]{rcr}Q_{A,B}^{T}X&\leq&\delta(\sigma^{EOR}_{k})\\ Q_{B,A}^{T}X&\leq&\delta(\sigma^{EOR}_{k})\\ \vdots\end{array}\right.$
@@
(Dual) $\min_{\lambda\geq 0}$ $\quad g(\lambda)=\delta(\sigma^{EOR}_{k})\overbrace{\sum_{\{A,B\}}(\lambda_{A,B}+\lambda_{B,A})}^{G(G-1)/2}+k\lambda_{k}+\sum_{i=1}^{n}\lambda_{i}^{\prime}$
(37) s.t. $\sum_{\{A,B\}}Q_{A,B}(\lambda_{A,B}-\lambda_{B,A})+\lambda_{k}+\lambda^{\prime}\geq\frac{P}{\sum_{g}nRel(g)}$
@@
We construct $\lambda^{\prime}_{i}$ corresponding to constraint (36) and $\lambda_{k}$ corresponding to constraint (select up to k elements) below.
@@
$\phi=\frac{2}{(G-1)\sum_{g=1}^{G}nRel(g)}\left(\sum_{\{A,B\}}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|\right)$ and $\delta_{max}=\max_{g}\left\{\frac{\sigma^{PRP,g}[1]}{nRel(g)}\right\}$ from Lemma D.4. ∎
@@
This lemma follows directly from the EOR ranking principle of choosing the candidate that minimizes $\delta(\sigma^{EOR}_{k})$ defined according to Eq. (8).
@@
For every pair of $\lambda_{A,g}$ and $\lambda_{g,A}$, where $g\in\{1,\cdots G\}$ and $g\neq A$, only one of $\lambda_{A,g},\lambda_{g,A}$ is $\geq 0$. Each of the $G-1$ terms inside the summation in Eq. (42) reduces to the two group case as follows. For $i>k_{A}$ and each {$A,g$}, the term evaluates to $\leq 0$ using Lemma C.2 and thus $\lambda_{i}^{\prime}$ is clipped to 00. Similarly, for $i\leq k_{A}$ and each {$A,g$} the term evaluates to $\geq 0$ and thus $\lambda_{i}^{\prime}\geq 0$.
@@
We now show that $\lambda_{k}\geq 0$. From Eq. (40),
@@
Each of the $G-1$ terms inside the summation in Eq. (43) reduces to the two group case. For each {$A,g$}, the term evaluates to $\geq 0$ using Lemma C.2 and thus $\lambda_{k}\geq 0$.
@@
Using the fact that $\lambda_{i}^{\prime}=0$ for $i>k_{A}$ from Lemma D.2, and substituting $\lambda_{k}$ from Eq. (40), we get
@@
Each of the $G-1$ terms inside the summation in Eq. (45) reduces to the two group case. For each {$A,g$}, the term evaluates to $\frac{p_{i}}{(G-1)\sum_{g}nRel(g)}$ using Lemma C.3 and thus the corresponding duality constraint is satisfied.
@@
Using the fact that $\lambda_{i}^{\prime}\geq 0$ for $i\leq k_{A}$ from Lemma D.2, and substituting $\lambda_{k}$ from Eq. (40), $\lambda_{i}^{\prime}$ for $i\leq k_{A}$ in (41), we get
@@
We argue by an inductive argument similar to the proof of Theorem 5.2. Consider the base case of $k=1$, when the first element is to be selected. The EOR algorithm will select according to Eq. (8) resulting in the lower $\delta(\sigma_{k=1})$. Thus, $\delta(\sigma_{k=1})$ is clearly $\leq\delta_{max}$.
@@
At the next prefix $k$, if the group that is selected has $\frac{nRel(.|\sigma_{k})}{nRel(.)}\leq\frac{nRel(g|\sigma_{k-1})}{nRel(g)}$, then $\delta(\sigma_{k})\leq\delta_{max}$. Note that $\delta(\sigma_{k})$ is always non-negative by definition from Eq. (8).
@@
Eq. (D.1) holds since group $1$ is now the group with maximum relevance proportion after adding $p_{i}$ the top most current element from group $1$. Group $2$ becomes the group with minimum relevance proportion.
@@
We have shown above that if the group with lowest relevance proportion at prefix $k-1$ (group $1$ in this case) is selected and its relevance proportion now exceeds the group with the highest relevance proportion at prefix $k-1$ (group $g$ in the case above), then $\delta(\sigma_{k})\leq\delta_{max}$. Thus, we can say that at least one group exists that satisfies $\delta_{max}$ EOR constraint at prefix $k$. This completes the proof that the EOR algorithm always provides a feasible ranking that satisfies $\delta_{max}=\max_{g\in\{1\cdots G\}}\left\{\frac{\sigma^{PRP,g}[1]}{nRel(g)}\right\}$ for $G$ groups. ∎
@@ E.1. Baselines
Thompson Sampling Ranking Policy ($\pi^{TS}$) (Singh et al., 2021)
For $\pi^{TS}$, binary relevances are drawn according to $r_{i}\sim\mathbb{P}(r_{i}|\mathcal{D})$, and candidates are sorted in decreasing order of relevance $r_{i}$ with their ranking randomized for the same value of relevance $r_{i}$.
@@
Demographic Parity ($\pi^{DP}$)
Candidates in each group are sorted in decreasing order of $\mathbb{P}(r_{i}|\mathcal{D})$ and selected such that the following constraint is minimized. This constraint is similar to the statistical parity variations introduced in (Yang and Stoyanovich, 2016).
@@
where $S(.)$ represents the size of the group. For a fair comparison with $\sigma^{EOR}$, we use Algorithm 1 and instead of minimizing Eq. (6), we minimize the above demographic parity constraint (49). We now discuss other variations of proportional representation constraints that have been introduced in prior literature (Celis et al., 2017, 2020, 2021). Generally, these constraints require that the disadvantaged group selected is at least a specific proportion $\alpha$ of top k.
@@
where $\alpha=\frac{S(B)}{S(A)+S(B)}$ and Eq. (50) is used as the fairness constraint while maximizing the utility to the principal. This type of representational constraint by definition requires the designation of a disadvantaged group. By designating B as the disadvantaged group, the constraint for proportional Rooney-Rule policy (Shen et al., 2023), which we denote by $\pi^{PRR}$ is as follows
@@
Figure 9. EOR criterion $\delta(\sigma_{k})$, costs of the ranking policies, and DCG Utility for Synthetic dataset with proportional Rooney-Rule like constraint, $\pi^{PRR}$. For group A we draw $S(A)=30$ relevance probabilities from $Powerlaw(\eta=5)$, and then draw for group B from $Powerlaw(\eta=0.5)$ until $nRel(A)\approx nRel(B)$.
@@
This criterion is anchored on the principle that a top-k ranking is fair when the proportion of disadvantaged candidates selected doesn’t fall far below a required minimum proportion $p$. This is formalized with a Binomial distribution, and a confidence level ($1-\alpha$). A function of the binomial cdf is computed apriori and is used as an input in the FAIR Algorithm. Since Binomial(p=0.5,n) corresponds to a ranking where at each position, a candidate from either group is selected randomly, FAIR is a ”softened” version of demographic parity (DP). As a result, FAIR is fundamentally different from Axiom 1 and Definition 4.1 derived from the uniform lottery fairness because, unlike DP, the uniform lottery is anchored on selecting an equal fraction of relevance from each group. Unlike $\pi^{EOR}$, $\pi^{FS}$ is oblivious to the relevance distribution and thus cannot take disparate uncertainty into account. FAIR also requires the normative designation of a disadvantaged group.
@@
For $\pi^{RA}$, we modify the baseline for fair rank aggregation in (Cachel and Rundensteiner, 2023) as follows. In fair rank aggregation, all $n$ candidates are ranked by $m$ voters to achieve a ranking with maximum consensus accuracy, where consensus may be according to different aggregation methods while achieving fairness of exposure w.r.t groups. (Cachel and Rundensteiner, 2023) proposes an algorithm that finds the consensus maximizing ranking and then swaps the candidates such that the equality of exposure is satisfied in that ranking. To adapt this baseline, we use the ranking from utility maximizing $\pi^{PRP}$ as the consensus ranking and use the algorithm from (Cachel and Rundensteiner, 2023) to swap elements in PRP ranking until the exposure constraint below is satisfied,
@@
A threshold of $0.95$ is used in experiments and on average over 100 runs, an exposure of $0.96\pm{0.01},0.96\pm{0.00},0.97\pm{0.00}$ is achieved for high, medium, and low levels of disparate uncertainty respectively in Table 1.